The equation for the instantaneous voltage across a discharging capacitor is given by 𝑣 = 𝑉𝑂𝑒 βˆ’ 𝑑 𝜏 , where 𝑉𝑂 is the initial voltage and 𝜏 is the time constant of the circuit. The tasks are to: a) Draw a graph of voltage against time for 𝑉𝑂 = 12𝑉 and 𝜏 = 2𝑠, between 𝑑 = 0𝑠 and 𝑑 = 10𝑠. b) Calculate the gradient at 𝑑 = 2𝑠 and 𝑑 = 4𝑠. c) Differentiate 𝑣 = 12𝑒 βˆ’ 𝑑 2 and calculate the value of 𝑑𝑣 𝑑𝑑 at 𝑑 = 2𝑠 and 𝑑 = 4𝑠. d) Compare your answers for part b and part c. e) Calculate the second derivative of the instantaneous voltage ( 𝑑 2 𝑣 𝑑𝑑2 ).


1
Expert's answer
2022-03-14T18:35:10-0400

Solution


The voltage across a discharging capacitor is given by

"v = {V_0}\,e - \tau t\"

Using


"{V_0}=12 \ V" and "\tau = 2\ s"


We have


"v = 12\,e - 2 t\"


This is a linear equation and its plot is shown below (graph of "v = {V_0}\,e - \tau t\" voltage against time for "{V_0}=12 \ V" and "\tau = 2\ s" , between "𝑡 = 0 \ 𝑠" and "𝑡 = 10\ 𝑠" .





Solution (b)


The above plot is a straight line. It means it has a constant gradient. From the equation "v = 12\,e - 2 t\" , the gradient is "m=-2".


Therefore, the gradient at 𝑑 = 2𝑠 is "m=-2" and 𝑑 = 4𝑠 is "m=-2".



Solution (c)


"v = 12\,e - 2 t\"


Its derivative is


"\frac{dv}{dt} = -2"


Therefore, "\frac{dv}{dt}" at 𝑑 = 2𝑠 and 𝑑 = 4𝑠, is "\frac{dv}{dt} = -2"



Solution (d)


From part (b) and part (c), we can say that "\frac{dv}{dt}" gives the gradient of the curve/ line "v = 12\,e - 2 t\" , which is constant "m=-2" for all values of "t" .



Solution (e)


"v = 12\,e - 2 t\"


Its first derivative is


"\frac{dv}{dt} = -2"


And the second derivative is


"\frac{{{d^2}v}}{{d{t^2}}} = 0\"